OpenStudy (anonymous):

Conversion: 72000 cm/hr^2 = ? m/s^2 Conversion: 500 km/minute = ? m/s

OpenStudy (baldymcgee6):

\[72000 \frac{ cm }{ hr^2 } * \frac{ 1m }{ 100cm } * (\frac{ 1 hr }{ 3600s })^2 = ??\]

OpenStudy (baldymcgee6):

Does that make sense to you?

OpenStudy (baldymcgee6):

\[72000 \frac{\cancel{cm} }{ \cancel{hr^2} } * \frac{ 1 m }{ 100 \cancel{cm} } * (\frac{ 1\cancel{ hr} }{ 3600s })^2 = ??\]

OpenStudy (baldymcgee6):

Look at the units how they all cancel each other nicely. That is how you can do all unit conversion problems!

OpenStudy (baldymcgee6):

@Sda370

OpenStudy (baldymcgee6):

\[500 \frac{ km }{ \min } * \frac{ 1000 m }{ 1 km } * \frac{ 1 \min }{ 60s } = ??\]

OpenStudy (anonymous):

Thank you, this was the way I learned it but didn't understand how to use it on these problems. You just won yourself a medal! =D

OpenStudy (baldymcgee6):

do you understand it now?

OpenStudy (anonymous):

Still writing it down, so the hour crosses out with the hour squared?

OpenStudy (baldymcgee6):

all your units should cancel except the ones you want left over, in this case m/s^2

OpenStudy (anonymous):

For the first question, is the answer 0.2 seconds squared?

OpenStudy (anonymous):

And the second, 8333.3m/s?

OpenStudy (baldymcgee6):

\[0.200 \frac{ m }{ s^2 }\]

OpenStudy (baldymcgee6):

\[8.33 * 10^3 m/s\]